OpenStudy (mani2596):

3 individuals take the elevator in 15-story building. What are the odds they all will get off on the same floor? Notice: they are all strangers to one another

OpenStudy (welshfella):

can we assume that they all start from the ground floor?

OpenStudy (mani2596):

yes

OpenStudy (welshfella):

so one person can go to any one of 14 floors. probability he gets out at floor 1 = 1/14 probability that the other 2 get out at this floor also = 1/14 * 1/14 * 1/14 Statistics is not my strong point so i would advise another opinion

OpenStudy (welshfella):

the other possiblities are they all get out at floors 2 to 14. So the required probability = 14 ( 1/14 * 1/14* 1/14) = 14 */ 2744 = 1 / 196

ganeshie8 (ganeshie8):

Looks neat to me...

OpenStudy (welshfella):

Hi ganeshie8 sorry gotta go right now